A problem in calculus is given to two students A and B whose chances of solving it are(1/3) and(1/4)
respectively. Find the probability of the problem being solved if both of them try
independently

Grade:Upto college level

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
As they are solving independently their probabilities of solving the problem will get multiplied.so P=(1/3)(1/4)=1/12.
Thanks and regards
Shaik Aasif
askIItians faculty